OpenStudy (anonymous):

How many real-number solutions does the equation have? –7x2 + 6x + 3 = 0 (1 point) one solution two solutions no solutions infinitely many solutions

OpenStudy (anonymous):

a quadratic equation can have at most 2 solutions. to check you need to calculate discriminant: D=b^2-4ac in your case: D=36-4.-7.3 D=36+4.7.3>0 so your equation has 2 solutions. if D=0 it would have had 1 if D < 0 none real solutions i am talking about so your answer is 2

OpenStudy (anonymous):

thank you. I was really confused on how to even start it. Thanks for the explanation.